OpenStudy (anonymous):

steps please : integrate 1/(2x+3) from 0 to 3

OpenStudy (anonymous):

Do a \(u=2x=3\) substitution.

OpenStudy (anonymous):

i know that the integral of 1/(2x+3) is 1/2 ln(2x+3)+C but i need the steps to do from ) to 3

OpenStudy (anonymous):

\(u = 2x+3\)

OpenStudy (anonymous):

Did you learn about u substitution yet?

OpenStudy (dumbcow):

u substitution: u = 2x+3 du = 2 dx \[\rightarrow \frac{1}{2}\int\limits_{0}^{3} \frac{1}{u} du\]

OpenStudy (anonymous):

yeah i know how to find the integral just forgot how to do from 0 to 3

OpenStudy (anonymous):

is it like: 1/2 ln (9) - 1/2 ln(-3)?

OpenStudy (dumbcow):

oh you need to use Fundamental Thm of Calc \[\int\limits_{a}^{b}f(x) dx = F(b) - F(a)\] so \[\frac{1}{2}[\ln (2*3+1) - \ln (2*0+1)]\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

wait shouldnt it be 1/2 (ln 2*3+3) and not 1?

hartnn (hartnn):

yup, its +3 and not +1 that was probably just a typo

OpenStudy (anonymous):

alright, was confused. Thanks!